### v1

# Manifold-aware Representation Learning for Degradation-agnostic Image Restoration

Abstract
Image Restoration (IR) aims to recover high-quality images from degraded inputs affected by various corruptions such as noise, blur, haze, rain, and low-light conditions. Despite recent advances, most existing approaches treat IR as a direct mapping problem, relying on shared representations across degradation types without modeling their structural diversity. In this work, we present MIRAGE, a unified and lightweight framework for all-in-one IR that explicitly decomposes the input feature space into three semantically aligned parallel branches, each processed by a specialized module—attention for global context, convolution for local textures, and MLP for channel-wise statistics. This modular decomposition significantly improves generalization and efficiency across diverse degradations. Furthermore, we introduce a cross-layer contrastive learning scheme that aligns shallow and latent features to enhance the discriminability of shared representations. To better capture the underlying geometry of feature representations, we perform contrastive learning in a Symmetric Positive Definite (SPD) manifold space rather than the conventional Euclidean space. Extensive experiments show that MIRAGE not only achieves new state-of-the-art performance across a variety of degradation types but also offers a scalable solution for real-world IR scenarios. Our code and models will be publicly available at our Project Page.
Figure 1: (a)-(d): Visual comparison for Denoising, Deraining, Composited Degradations (low-light, haze, and snow), and underwater image enhancement. (e): The average PSNR and SSIM comparison across 4 challenging all-in-one and 1 zero-shot settings (Please zoom in for a better view).
1 Introduction
Image Restoration (IR) is a fundamental yet challenging problem in computer vision, aiming to reconstruct clean images from their degraded versions affected by a variety of real-world corruptions, including noise, blur, haze, rain, low-light conditions, and more zamir2022restormer ; li2023efficient ; ren2024sharing ; zamfir2025moce ; ma2024shapesplat ; potlapalli2023promptir . While recent advances have made great strides by designing powerful architectures based on Convolutional Neural Networks (CNNs), Multi-Layer Perceptrons (MLPs), and Transformers dosovitskiy2020image ; ren2024sharing ; zhu2023good , most existing methods view restoration as a direct mapping problem—learning a global function that transforms the corrupted input into its clean counterpart. These approaches typically rely on shared representations across all degradation types, often overlooking the structural distinctions among them. As a result, they struggle to generalize when confronted with unseen or mixed degradations.
In reality, different degradation types, such as noise, haze, rain, snow, and motion blur, arise from fundamentally different physical processes and can be roughly categorized as additive, multiplicative, or kernel-based corruptions. Simultaneously, basic architectural modules exhibit distinct processing biases: convolutional filters excel at modeling local texture, attention mechanisms capture global dependencies, and MLPs often enhance channel-wise statistics. This leads to a key insight: To handle diverse and complex degradations effectively, a restoration model should be equipped with multiple processing capabilities. However, designing such a unified framework is non-trivial. A naive combination of CNNs, MLPs, and attention units often leads to redundancy, optimization difficulty, and computational inefficiency. In this work, we revisit the widely observed redundancy in attention mechanisms, particularly along the channel dimension venkataramanan2023skip ; ji2025always ; dong2021attention . This phenomenon has been extensively validated in both the natural language processing (NLP) and vision communities, showing that many attention heads or channels contribute minimally to performance. Rather than avoiding this redundancy, we propose to exploit it to build an efficient, modular, and generalizable architecture. We present MIRAGE, a unified and lightweight IR framework that explicitly decomposes feature processing into three semantically aligned branches. Specifically, we split the input feature map along the channel dimension and assign each partition to a specialized module: attention for global context modeling, convolution for local texture enhancement, and MLP for channel-wise refinement. This structured decomposition not only enhances expressiveness but also reduces computational complexity, yielding a powerful backbone that performs competitively across a wide range of degradation scenarios.
While MIRAGE already provides strong results and surpasses many state-of-the-art methods, we observe that its performance primarily stems from its architectural expressiveness rather than an explicit understanding of degradation-specific semantics. Motivated by recent efforts such as AirNet and PromptIR, which incorporate degradation cues into the learning process, we argue that degradation-guided processing can make the shared representation more discriminative and significantly improve generalization across diverse restoration tasks. To this end, we introduce a novel cross-layer contrastive learning strategy. Unlike conventional contrastive learning, which relies on complex augmentations, heavy architectures, or multi-stage training, we take a simple yet effective approach. Inspired by deeply supervised networks lee2015deeply , we hypothesize that natural contrastive pairs exist within the model itself, i.e., between shallow and deeper latent representations. These layers carry complementary characteristics: shallow features are rich in spatial details and noise sensitivity, while latent features are more abstract and semantically consistent. Aligning them can improve both robustness and generalization. Furthermore, instead of applying contrastive loss in Euclidean space, which can misrepresent feature similarity, we perform it in the Symmetric Positive Definite (SPD) manifold space, leading to more meaningful geometric alignment.
Inspired by the metaphor of uncovering the latent clean scene beneath complex degradations, just as a mirage reveals a hidden reality, our method learns a degradation-agnostic representation by dynamically balancing global context, local structure, and channel-wise distribution. This synergy between architectural modularity and contrastive regularization makes MIRAGE a powerful and general-purpose backbone for challenging all-in-one image restoration.
Our main contributions are as follows:
  • We propose MIRAGE, a lightweight and unified IR framework that synergistically integrates attention, convolution, and MLP through structured channel-wise decomposition, enabling robust performance across diverse degradation types.
  • We introduce a simple yet effective cross-layer contrastive learning strategy that aligns shallow and latent features within the SPD manifold space, thereby enhancing feature discriminability and generalization. Our approach respects the intrinsic geometry of deep representations and improves restoration quality, all without relying on data augmentation or multi-stage training.
  • Extensive experiments across diverse restoration tasks demonstrate the effectiveness and efficiency of our approach. We hope MIRAGE can serve as a strong and practical baseline to inspire future research in all-in-one image restoration.
2 Related Work
Image Restoration (IR) with Various Basic Architectures. IR aims to address a highly ill-posed problem: reconstructing high-quality images from their degraded versions. Given its broad importance, IR has been widely applied in numerous applications richardson1972bayesian ; xie2025mat ; banham1997digital ; li2023lsdir ; zamfir2024details . Early IR methods primarily relied on model-based solutions that searched for closed-form solutions to predefined formulations. With the advent of deep neural networks, learning-based approaches have rapidly gained popularity. A wide range of methods have emerged, including regression-based techniques lim2017enhanced ; lai2017deep ; liang2021swinir ; chen2021learning ; li2023efficient ; zhang2024transcending and generative model-based pipelines gao2023implicit ; wang2022zero ; luo2023image ; yue2023resshift ; zhao2024denoising , built upon convolutional dong2015compression ; zhang2017learning ; zhang2017beyond ; wang2018recovering , MLP-based tu2022maxim , state space models guo2024mambair ; zhu2024vision ; mamba ; mamba2 , and Vision Transformer (ViT)-based architectures liang2021swinir ; ren2023masked ; li2023efficient ; zamir2022restormer ; dosovitskiy2020image ; liu2023spatio . Despite the impressive performance of recent state-of-the-art methods, most IR solutions are still designed to address specific degradation types, such as denoising zhang2017learning ; zhang2019residual , dehazing ren2020single ; wu2021contrastive , deraining jiang2020multi ; ren2019progressive , deblurring kong2023efficient ; ren2023multiscale , and others.
Degradation-agnostic Image Restoration. While training task-specific models to handle individual types of degradation can be effective, it poses practical limitations due to the need for separate models for each degradation. In real-world scenarios, images often suffer from a mixture of degradations and artifacts, making it difficult to address each type independently. Task-specific solutions also demand considerable computational and storage resources, significantly increasing their environmental footprint. To overcome these limitations, the emerging field of All-in-One image restoration focuses on single-blind models capable of handling multiple degradation types simultaneously zamfir2025moce ; zeng2025vision ; zheng2024learning . For example, AirNet li2022all achieves blind All-in-One image restoration by using contrastive learning to derive degradation representations from corrupted images, which are then leveraged to reconstruct clean images. Building on this, IDR zhang2023ingredient tackles the problem by decomposing degradations into fundamental physical components and applying a two-stage meta-learning strategy. More recently, the prompt-based paradigm potlapalli2023promptir ; wang2023promptrestorer ; li2023prompt has introduced a visual prompt learning module, enabling a single model to better handle diverse degradation types by leveraging the discriminative capacity of learned visual prompts. Extending this idea, some works further model prompts from a frequency perspective cui2025adair or propose more complex architectures with additional datasets dudhane2024dynamic . However, visual prompt modules often result in increased training time and decreased efficiency cui2025adair . In contrast, our work aims to improve the model’s ability to capture representative degradation cues without relying on heavy or complex prompt designs. Our goal is to develop an All-in-One image restorer that remains both computationally efficient and environmentally sustainable.
3 Preliminary: Degradation-Aware Architectures for Image Restoration
Image Degradation and Restoration. Image restoration aims to recover a clean image $\mathbf{x}$ from a degraded observation $\mathbf{y}$, commonly modeled as:
$\mathbf{y}=\mathcal{D}(\mathbf{x})+\mathbf{n},$ (1)
where $\mathcal{D}(\cdot)$ is a degradation operator and $\mathbf{n}$ is noise. Real-world degradations are diverse in nature—ranging from additive (e.g., Gaussian noise, rain: $\mathbf{y}=\mathbf{x}+\mathbf{n}$), multiplicative (e.g., haze, speckle: $\mathbf{y}=\mathbf{x}\cdot\mathbf{m}$), to convolutional degradations (e.g., blur, super-resolution: $\mathbf{y}=\mathbf{k}*\mathbf{x}+\mathbf{n}$).
Such degradations are often entangled and spatially variant, forming compound pipelines:
$\mathbf{y}=\mathcal{D}_{3}\big{(}\mathcal{D}_{2}(\mathcal{D}_{1}(\mathbf{x}))\big{)}+\mathbf{n},$ (2)
where each $\mathcal{D}_{i}$ captures a distinct degradation type. Addressing this diversity requires models capable of both local detail preservation and global structural reasoning.
Architectural Biases for Degradation Modeling. To this end, modern deep networks adopt different inductive biases: CNNs capture local spatial patterns through convolution: $\mathbf{y}_{p}=\sum_{i\in\mathcal{N}(p)}w_{i}\cdot\mathbf{x}_{i}$. Their locality makes them effective for uniform or spatially invariant degradations. Transformers exploit global self-attention: $\mathbf{y}_{i}=\sum_{j}\alpha_{ij}\cdot\mathbf{V}_{j}$, enabling modeling of non-uniform, structured degradations such as haze or patterned noise. MLPs, especially token-mixing variants, use position-wise transformations: $\mathbf{y}=\mathbf{W}_{2}\cdot\phi(\mathbf{W}_{1}\cdot\mathbf{x})$, offering flexibility with weak spatial priors.
Each paradigm exhibits strengths and limitations in handling different degradation types. CNNs excel in local fidelity, Transformers in global context, and MLPs offer flexible feature interactions but lack inherent spatial priors. When applied to vision tasks, they often require large parameter counts and are less effective alone for structured degradations. These complementary traits motivate unified architectures that integrate them for robust, degradation-aware restoration in the wild.
4 The Proposed MIRAGE
Figure 2: (a)-(c): The most adopted all-in-one image restoration encodedr-decoder pipelines. (d): The toy illustration of our SPD contrastive pipeline. (e): The overall framework of the proposed MIRAGE : i.e., a convolutional patch embedding, a U-shape encoder-decoder main body, an extra refined block, and the proposed SPD contrastive learning algorithm. (f): Structure of each mixed degradation adaptation block (MDAB).
Overview. Unlike previous methods that adopt a one-to-one mapping strategy for multiple degradations as illustrated in Fig. 2(a), which requires training a separate model for each type of degradation, our approach is more unified. Also, unlike architectures that use multiple encoders mapping to a single decoder, as shown in Fig. 2(b), which significantly increases the model size and complexity, we pursue a more efficient design. Moreover, in contrast to prompt-based methods that rely on large-scale models or introduce visual/textual prompts (Fig. 2(c)), we propose a simple yet effective mixed-backbone architecture (Fig. 2(d)). This backbone serves as a strong restoration model (Sec. 4.1), and its performance is further enhanced through cross-layer contrastive learning in the SPD space between shallow and latent features (Sec. 4.2).
4.1 Mixed Degradation Adaptation Block for AnyIR
Redundancy in MHAs Opens Opportunities for Hybrid Architectures. Redundancy has long been recognized as a fundamental limitation in multi-head self-attention (MHA), the core building block of Transformers, in both NLP and vision domains nguyen2022improving ; nguyen2022improvingkeys ; xiao2024improving ; brodermann2025cafuser ; wang2022improved ; venkataramanan2023skip . Prior studies have shown that not all attention heads contribute equally, i.e., some are specialized and crucial, while others can be pruned with negligible impact. This inherently implies redundancy in the channel dimension, as MHA outputs are concatenated along this axis. To empirically verify this redundancy in the context of image restoration (IR), we analyze intermediate features from a lightweight attention-only model (details in the Appendix).
Figure 3: Channel redundancy analysis across multiple feature scales. (a) Cumulative explained variance curves from PCA applied to the channel dimension of features from 1-4 scales and one latent scale. (b) Normalized singular value spectra (in log scale) of the same features via SVD. Latent feature in both plots means the channel-wise projected 4th Scale feature.
Specifically, we compute the cumulative explained variance via PCA and the normalized singular value spectra via SVD across multiple feature scales, as shown in Fig.3. As illustrated in Fig. 3(a), earlier scales (e.g., 1st Scale) require significantly fewer principal components to retain most of the variance, suggesting high redundancy. Fig. 3(b) further supports this observation, with a sharper singular value decay at shallower stages, indicating stronger low-rank structure in channel-wise representations. Even at the deepest stage (e.g., 4th Scale), achieving 90% variance only requires 28 out of 192 components ($\approx 19\%$), confirming that redundancy persists throughout.
This insight motivates a departure from traditional head/channel pruning. Instead of discarding redundant capacity, we propose to repurpose it by splitting the channel dimension into three parts and feeding them into distinct architectural branches, e.g., attention, convolution, and MLP. This hybrid formulation leverages complementary inductive biases and makes full use of available representational space, offering a principled and efficient alternative to the previous pure MSA-based designs.
Algorithm 1 Mixed Parallel Degradation Adaptation
1:$F_{\text{in}}^{\text{att}},\ F_{\text{in}}^{\text{conv}},\ F_{\text{in}}^{\text{mlp}}$triangleright Input features from three branches
2:$F_{\text{out}}$triangleright Final fused output
3:[Att] Attention Path
4:$Q,K,V\leftarrow\text{Linear}(F_{\text{in}}^{\text{att}})$triangleright Projection to attention tokens
5:$F_{\text{out}}^{\text{att}}\leftarrow\texttt{Softmax}(\frac{QK^{\top}}{\sqrt{d}})V$triangleright Multi-head self-attention
6:[Conv] Dynamic Convolution Path
7:$F^{\prime}\leftarrow\text{Conv1x1}(\text{Norm}(F_{\text{in}}^{\text{conv}}))$triangleright Normalization and expansion
8:$\gamma,\beta,\alpha\leftarrow\text{Split}(F^{\prime})$triangleright Gating, intermediate, convolutional paths
9:$\alpha^{\prime}\leftarrow\text{DynamicDepthwiseConv}(\alpha)$triangleright Content-adaptive depthwise conv
10:$\hat{F}\leftarrow\sigma(\gamma/\tau)\cdot\text{Concat}(\beta,\alpha^{\prime})$triangleright Gated local enhancement
11:$F_{\text{out}}^{\text{conv}}\leftarrow\text{Conv1x1}(\hat{F})+F_{\text{in}}^{\text{conv}}$triangleright Residual projection
12:[MLP] MLP Path
13:$F_{\text{out}}^{\text{mlp}}\leftarrow\text{MLP}(F_{\text{in}}^{\text{mlp}})$triangleright Channel-wise transformation brings more non-linearity
14:[Fusion] Inter-Branch Mutual Fusion
15:$F_{\text{out}}^{\text{att}^{\prime}}\leftarrow F_{\text{out}}^{\text{att}}+\lambda_{\text{att}}\cdot\sigma(F_{\text{out}}^{\text{conv}}+F_{\text{out}}^{\text{mlp}})$triangleright Fuse conv and MLP into attention
16:$F_{\text{out}}^{\text{conv}^{\prime}}\leftarrow F_{\text{out}}^{\text{conv}}+\lambda_{\text{conv}}\cdot\sigma(F_{\text{out}}^{\text{att}}+F_{\text{out}}^{\text{mlp}})$triangleright Fuse attention and MLP into conv
17:$F_{\text{out}}^{\text{mlp}^{\prime}}\leftarrow F_{\text{out}}^{\text{mlp}}+\lambda_{\text{mlp}}\cdot\sigma(F_{\text{out}}^{\text{att}}+F_{\text{out}}^{\text{conv}})$triangleright Fuse attention and conv into MLP
18:Output Projection
19:$F_{\text{out}}^{\text{fuse}}\leftarrow\text{Project}(\text{Concat}(F_{\text{out}}^{\text{att}^{\prime}},\ F_{\text{out}}^{\text{conv}^{\prime}},\ F_{\text{out}}^{\text{mlp}^{\prime}}))$triangleright Final unified representation
20:return $F_{\text{out}}^{\text{fuse}}$
Parallel Design Brings More Efficiency. The preceding redundancy analysis reveals that many channels across scales encode overlapping information. Rather than discarding this capacity via pruning, we strategically re-purpose it through a structurally parallel design. As illustrated in Fig. 2(f), the input feature $F_{\text{in}}\in\mathbb{R}^{h\times w\times c}$ is evenly divided along the channel dimension into three sub-tensors: $F_{\text{in}}^{\text{att}}$, $F_{\text{in}}^{\text{conv}}$, and $F_{\text{in}}^{\text{mlp}}$, corresponding to attention, convolution, and MLP branches. Each branch independently processes its input using lightweight modules that specialize in distinct inductive biases, i.e., global dependency modeling, local pattern extraction, and channel-wise transformation. This design reduces computation (each operation handles only a fraction of the channels) while enhancing representational diversity through architectural heterogeneity (see Lines 1-8 Alg. 1).
Inter-Branch Mutual Fusion Injects Expressivity Before FFN. While the parallel design improves efficiency and modularity, it inevitably reduces interaction across branches. To mitigate this, Lines 9–13 of Alg. 1 introduce an inter-branch mutual fusion mechanism, where each branch is enhanced via gated aggregation of the other two, modulated by learnable coefficients $\lambda$. This introduces non-linearity and cross-path context blending, reinforcing feature complementarity before unification. The fused representation is more expressive, forming a compact and effective pre-FFN encoder.
Compared to standard attention-only models where $F_{\text{out}}^{\text{att}}$ is essentially a linear combination modulated by Softmax weights, the fused output in Alg. 1 introduces richer non-linear interactions. This enhances the model’s ability to fit complex degradation mappings, making it more suitable for real-world scenarios with mixed or ambiguous degradations. Subsequently, layer normalization, a feed-forward network (FFN), and a residual connection are applied: $F_{\text{out}}=\operatorname{FFN}(\operatorname{Norm}(F_{\text{out}}^{\text{fuse}}))+F_{\text{out}}^{\text{fuse}}$. This sequence stabilizes feature distributions and further boosts expressiveness.
Figure 4: (a)-(d): The channel-wise similarity matrix from the 1st Scale ($H\times W\times C$) to the 4th Scale ($\nicefrac{{H}}{{8}}\times\nicefrac{{W}}{{8}}\times 8C$). (e): The channel-wise similarity matrix of (d) after channel-wise projection.
4.2 Shallow-Latent Contrastive Learning via SPD Manifold Alignment
Shallow-Latent Feature Pairs are Naturally Contrastive Pairs. Features extracted at different depths exhibit fundamentally different statistical properties. As shown in Fig. 4, shallow-stage features (e.g., Scale1) present sparse and decorrelated channel distributions, while deeper layers (e.g., Scale4) become increasingly redundant and concentrated. This trend is quantitatively supported by the effective rank ratio across scales, which increases from only 4.2% ($\nicefrac{{1}}{{24}}$ at 1st Scale) to 16.1% ($\nicefrac{{31}}{{192}}$ at 4th Scale). However, by compressing the deep features through a lightweight MLP, we obtain a latent representation with a significantly higher rank ratio of 36.5% ($\nicefrac{{35}}{{96}}$), indicating a more decorrelated and expressive embedding. This structural disparity between sparse, localized shallow features and compressed, semantic latent ones naturally defines a contrastive pairing without requiring additional augmentation. We leverage this depth-asymmetric contrast to impose consistency across stages, enabling better semantic alignment and stronger representational generalization under complex degradation conditions. Note that this case study is conducted under noise degradation; however, similar trends are consistently observed for other degradations as well. Please refer to our Supplementary Materials (i.e., Supp. Mat.) for more details.
SPD Manifold Space Contrastive Learning Leads to More Discriminative Representations. To enhance representation consistency across depth, we introduce a contrastive objective defined over SPD (Symmetric Positive Definite) manifold features. Given shallow features $F_{\text{shallow}}\in\mathbb{R}^{C_{s}\times H\times W}$ and latent features $F_{\text{latent}}\in\mathbb{R}^{C_{l}\times H^{\prime}\times W^{\prime}}$, we first reduce their channel dimensions via $1\times 1$ convolutions. The resulting tensors are reshaped into feature matrices $X_{s},X_{l}\in\mathbb{R}^{C\times N}$ with $N=H\times W$, and their second-order statistics are computed as:
$\mathbf{C}_{s}=\frac{1}{N-1}(X_{s}-\mu_{s})(X_{s}-\mu_{s})^{\top}+\epsilon\mathbf{I},\quad\mathbf{C}_{l}=\frac{1}{N^{\prime}-1}(X_{l}-\mu_{l})(X_{l}-\mu_{l})^{\top}+\epsilon\mathbf{I},$ (3)
where $\mu$ is the mean across spatial dimensions, and $\epsilon\mathbf{I}$ ensures numerical stability and positive definiteness. The SPD matrices $\mathbf{C}_{s},\mathbf{C}_{l}\in\mathbb{R}^{C\times C}$ are vectorized and projected to a contrastive embedding space via shallow MLPs:
$z_{s}=\operatorname{Norm}(W_{s}\cdot\text{vec}(\mathbf{C}_{s})),\quad z_{l}=\operatorname{Norm}(W_{l}\cdot\text{vec}(\mathbf{C}_{l})),$ (4)
where $W_{s},W_{l}$ are learnable projection layers, and $\operatorname{Norm}(\cdot)$ denotes $\ell_{2}$-normalization. We then apply an InfoNCE-style contrastive loss to align the shallow and latent embeddings:
$\mathcal{L}_{\text{SPD}}=-\log\frac{\exp\left(\operatorname{sim}(z_{s},z_{l})/\tau\right)}{\sum\limits_{z_{l}^{\prime}}\exp\left(\operatorname{sim}(z_{s},z_{l}^{\prime})/\tau\right)},$ (5)
where $\operatorname{sim}(\cdot,\cdot)$ denotes cosine similarity, and $\tau$ is a temperature parameter. Unlike conventional Euclidean contrastive learning which treats feature vectors as flat points our SPD-based method preserves second-order channel dependencies, enabling richer structural supervision. This manifold-aware regularization aligns local and semantic features across depth, enhances discriminability, and introduces no additional inference cost.
5 Experiments
We conduct experiments adhering to the protocols of prior general image restoration works potlapalli2023promptir ; zhang2023ingredient under four settings: (i) All-in-One (3Degradations), (ii) All-in-One (5Degradations), (iii) Mixed Degradation Setting, (iv) Adverse Weather Removal Setting, and (v) Zero-Shot Setting. More experimental details and the dataset introduction are provided in our Supp. Mat.
5.1 SOTA Comparison.
3 Degradations. We evaluate our all-in-one restorer, MIRAGE , against other specialized methods listed in Tab. 1, all trained on three degradations: dehazing, deraining, and denoising. MIRAGE consistently outperforms all the comparison methods, even for those with the assistance of language, multi-task, or prompts. Notably, even our 6M tiny model outperforms the baseline method PromptIR potlapalli2023promptir by 0.71dB on average. And the 10M small model achieves the best performance across all the metrics, while maintaining 60% fewer parameters compared to MoCE-IR zamfir2025moce .
Table 1: Comparison to state-of-the-art on three degradations. PSNR (dB, $\uparrow$) and SSIM ($\uparrow$) metrics are reported on the full RGB images. Best performances is highlighted. ‘-’ means unreported results.
Method Venue. Params. Dehazing Deraining Denoising Average
[…]
BRDNet tian2000brdnet Neural Networks’22 - 23.23 .895 27.42 .895 32.26 .898 29.76 .836 26.34 .693 27.80 .843
LPNet gao2019dynamic CVPR’19 - 20.84 .828 24.88 .784 26.47 .778 24.77 .748 21.26 .552 23.64 .738
FDGAN dong2020fdgan AAAI’20 - 24.71 .929 29.89 .933 30.25 .910 28.81 .868 26.43 .776 28.02 .883
DL fan2019dl TPAMI’19 2M 26.92 .931 32.62 .931 33.05 .914 30.41 .861 26.90 .740 29.98 .876
MPRNet zamir2021multi CVPR’21 16M 25.28 .955 33.57 .954 33.54 .927 30.89 .880 27.56 .779 30.17 .899
AirNet li2022all CVPR’22 9M 27.94 .962 34.90 .967 33.92 .933 31.26 .888 28.00 .797 31.20 .910
NDR yao2024neural TIP’24 28M 25.01 .860 28.62 .848 28.72 .826 27.88 .798 26.18 .720 25.01 .810
PromptIR potlapalli2023promptir NeurIPS’23 36M 30.58 .974 36.37 .972 33.98 .933 31.31 .888 28.06 .799 32.06 .913
MoCE-IR-S zamfir2025moce CVPR’25 11M 30.98 .979 38.22 .983 34.08 .933 31.42 .888 28.16 .798 32.57 .916
AdaIR cui2025adair ICLR’25 29M 31.06 .980 38.64 .983 34.12 .935 31.45 .892 28.19 .802 32.69 .918
MoCE-IR zamfir2025moce CVPR’25 25M 31.34 .979 38.57 .984 34.11 .932 31.45 .888 28.18 .800 32.73 .917
MIRAGE -T (Ours) 2025 6M 31.81 .982 38.44 .983 34.05 .935 31.40 .892 28.14 .802 32.77 .919
MIRAGE -S (Ours) 2025 10M 31.86 .981 38.94 .985 34.12 .935 31.46 .891 28.19 .803 32.91 .919
Methods with the assistance of vision language, multi-task learning, natural language prompts, or multi-modal control
DA-CLIP luo2023controlling ICLR’24 125M 29.46 .963 36.28 .968 30.02 .821 24.86 .585 22.29 .476 - -
ArtPromptIR wu2024harmony ACM MM’24 36M 30.83 .979 37.94 .982 34.06 .934 31.42 .891 28.14 .801 32.49 .917
InstructIR-3D conde2024high ECCV’24 16M 30.22 .959 37.98 .978 34.15 .933 31.52 .890 28.30 .804 32.43 .913
UniProcessor duan2025uniprocessor ECCV’24 1002M 31.66 .979 38.17 .982 34.08 .935 31.42 .891 28.17 .803 32.70 .918
VLU-Net zeng2025vision CVPR’25 35M 30.71 .980 38.93 .984 34.13 .935 31.48 .892 28.23 .804 32.70 .919
Table 2: Comparison to state-of-the-art on five degradations. PSNR (dB, $\uparrow$) and SSIM ($\uparrow$) metrics are reported on the full RGB images with $(^{\ast})$ denoting general image restorers, others are specialized all-in-one approaches. Best performance is highlighted.
Method Venue Params. Dehazing Deraining Denoising Deblurring Low-Light Average
[…]
NAFNet chen2022simple ECCV’22 17M 25.23 .939 35.56 .967 31.02 .883 26.53 .808 20.49 .809 27.76 .881
DGUNet mou2022deep CVPR’22 17M 24.78 .940 36.62 .971 31.10 .883 27.25 .837 21.87 .823 28.32 .891
SwinIR liang2021swinir ICCVW’21 1M 21.50 .891 30.78 .923 30.59 .868 24.52 .773 17.81 .723 25.04 .835
Restormer zamir2022restormer CVPR’22 26M 24.09 .927 34.81 .962 31.49 .884 27.22 .829 20.41 .806 27.60 .881
MambaIR guo2024mambair ECCV’24 27M 25.81 .944 36.55 .971 31.41 .884 28.61 .875 22.49 .832 28.97 .901
DL fan2019dl TPAMI’19 2M 20.54 .826 21.96 .762 23.09 .745 19.86 .672 19.83 .712 21.05 .743
Transweather valanarasu2022transweather CVPR’22 38M 21.32 .885 29.43 .905 29.00 .841 25.12 .757 21.21 .792 25.22 .836
TAPE liu2022tape ECCV’22 1M 22.16 .861 29.67 .904 30.18 .855 24.47 .763 18.97 .621 25.09 .801
AirNet li2022all CVPR’22 9M 21.04 .884 32.98 .951 30.91 .882 24.35 .781 18.18 .735 25.49 .847
IDR zhang2023ingredient CVPR’23 15M 25.24 .943 35.63 .965 31.60 .887 27.87 .846 21.34 .826 28.34 .893
PromptIR potlapalli2023promptir NeurIPS’23 36M 26.54 .949 36.37 .970 31.47 .886 28.71 .881 22.68 .832 29.15 .904
MoCE-IR-S zamfir2025moce CVPR’25 11M 31.33 .978 37.21 .978 31.25 .884 28.90 .877 21.68 .851 30.08 .913
AdaIR cui2025adair ICLR’25 29 30.53 .978 38.02 .981 31.35 .889 28.12 .858 23.00 .845 30.20 .910
MoCE-IR zamfir2025moce CVPR’25 25M 30.48 .974 38.04 .982 31.34 .887 30.05 .899 23.00 .852 30.58 .919
MIRAGE -T (Ours) 2025 6M 31.35 .979 38.24 .983 31.35 .891 27.98 .850 23.11 .854 30.41 .912
MIRAGE -S (Ours) 2025 10M 31.45 .980 38.92 .985 31.41 .892 28.10 .858 23.59 .858 30.68 .914
Methods with the assistance of natural language prompts or multi-task learning
InstructIR-5D conde2024high ECCV’24 16M 36.84 .973 27.10 .956 31.40 .887 29.40 .886 23.00 .836 29.55 .908
ArtPromptIR wu2024harmony ACM MM’24 36M 29.93 .908 22.09 .891 29.43 .843 25.61 .776 21.99 .811 25.81 .846
VLU-Net zeng2025vision CVPR’25 35M 30.84 .980 38.54 .982 31.43 .891 27.46 .840 22.29 .833 30.11 .905
5 Degradations. Extending the three degradation tasks to include deblurring and low-light enhancement li2022all ; zhang2023ingredient , we validate our method’s comprehensive performance in an All-in-One setting. As shown in Tab. 2, MIRAGE -S effectively leverages degradation-specific features, surpassing PromptIR potlapalli2023promptir , MoCE-IR-S zamfir2025moce , AdaIR cui2025adair , and VLU-Net zeng2025vision by an average of 1.53dB, 0.6dB, 0.48dB, and 0.57dB with lower parameters. Notably, our tiny model (6M) also achieves competitive results (a second-best average PSNR) compared to the MoCE-IR (25M) and outperforms all other comparison methods, even those with the assistance of other modalities, multi-task learning, or pretraining.
Composited Degradation Setting.
Table 3: Comparison to state-of-the-art on composited degradations. PSNR (dB, $\uparrow$) and SSIM ($\uparrow$) are reported on the full RGB images. Our method consistently outperforms even larger models, with favorable results in composited degradation scenarios.
Method Params. CDD11-Single CDD11-Double CDD11-Triple Avg.
Low (L) Haze (H) Rain (R) Snow (S) L+H L+R L+S H+R H+S L+H+R L+H+S
AirNet li2022all 9M 24.83 .778 24.21 .951 26.55 .891 26.79 .919 23.23 .779 22.82 .710 23.29 .723 22.21 .868 23.29 .901 21.80 .708 22.24 .725 23.75 .814
PromptIR potlapalli2023promptir 36M 26.32 .805 26.10 .969 31.56 .946 31.53 .960 24.49 .789 25.05 .771 24.51 .761 24.54 .924 23.70 .925 23.74 .752 23.33 .747 25.90 .850
WGWSNet ZhuWFYGDQH23 26M 24.39 .774 27.90 .982 33.15 .964 34.43 .973 24.27 .800 25.06 .772 24.60 .765 27.23 .955 27.65 .960 23.90 .772 23.97 .771 26.96 .863
WeatherDiff weather_diffusion 83M 23.58 .763 21.99 .904 24.85 .885 24.80 .888 21.83 .756 22.69 .730 22.12 .707 21.25 .868 21.99 .868 21.23 .716 21.04 .698 22.49 .799
OneRestore guo2024onerestore 6M 26.48 .826 32.52 .990 33.40 .964 34.31 .973 25.79 .822 25.58 .799 25.19 .789 29.99 .957 30.21 .964 24.78 .788 24.90 .791 28.47 .878
MoCE-IR zamfir2025moce 11M 27.26 .824 32.66 .990 34.31 .970 35.91 .980 26.24 .817 26.25 .800 26.04 .793 29.93 .964 30.19 .970 25.41 .789 25.39 .790 29.05 .881
MIRAGE (ours) 6M 27.13 .830 32.39 .989 34.23 .969 35.57 .978 26.04 .823 26.21 .807 26.07 .799 29.49 .962 29.72 .967 25.17 .793 25.41 .793 28.86 .883
MIRAGE (ours) 10M 27.41 .833 33.12 .992 34.66 .971 35.98 .981 26.55 .828 26.53 .810 26.33 .803 30.32 .965 30.27 .969 25.59 .801 25.86 .799 29.33 .887
To simulate more realistic restoration scenarios, we extend the setting from OneRestore guo2024onerestore by including not only diverse single degradations—rain, haze, snow, low illumination—but also composite degradations where multiple types are combined within the same image. This results in a total of eleven distinct restoration settings. As shown in Tab. 3, our Tiny (6M) and Small (10M) models outperform OneRestore guo2024onerestore (6M) by 0.39 dB and 0.86 dB on average, respectively. Compared to the recent state-of-the-art MoCE-IR zamfir2025moce (11M), our Small model still achieves 0.28 dB higher performance while being more compact (10M vs. 11M). These results demonstrate the effectiveness of our method, especially in handling complex, mixed degradations.
Table 4: Comparisons for 4-task adverse weather removal. Missing values are denoted by ’–’.
Method Venue Snow100K-S Snow100K_dataset Snow100K-L Snow100K_dataset Outdoor-Rain test1_dataset_allweather RainDrop raindrop Average
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
All-in-One as2020 CVPR’20 28.33 .882 24.71 .898 31.12 .927 28.05 .902
TransWeather Transweather CVPR’22 32.51 .934 29.31 .888 28.83 .900 30.17 .916 30.20 .909
Chen et al. ChenHTYDK22 CVPR’22 34.42 .947 30.22 .907 29.27 .915 31.81 .931 31.43 .925
WGWSNet ZhuWFYGDQH23 CVPR’23 34.31 .946 30.16 .901 29.32 .921 32.38 .938 31.54 .926
WeatherDiff64 weather_diffusion TPAMI’23 35.83 .957 30.09 .904 29.64 .931 30.71 .931 31.57 .931
WeatherDiff128 weather_diffusion TPAMI‘23 35.02 .952 29.58 .894 29.72 .922 29.66 .923 31.00 .923
AWRCP AWRCP_YeCBSXJYCL23 ICCV’23 36.92 .965 31.92 .934 31.39 .933 31.93 .931 33.04 .941
GridFormer Gridformer IJCV’24 37.46 .964 31.71 .923 31.87 .933 32.39 .936 33.36 .939
MPerceiver AiHZW024 CVPR’24 36.23 .957 31.02 .916 31.25 .925 33.21 .929 32.93 .932
DTPM DTPM_0001CCXQL024 CVPR’24 37.01 .966 30.92 .917 30.99 .934 32.72 .944 32.91 .940
Histoformer Histoformer_SunRGWC24 ECCV’24 37.41 .966 32.16 .926 32.08 .939 33.06 .944 33.68 .944
MIRAGE -S (Ours) 2025 37.97 .973 32.33 .929 32.82 .949 32.78 .945 33.98 .949
Adverse Weather Removal Setting. Following valanarasu2022transweather ; zhu2023learning , We test our MIRAGE on three challenging deweathering tasks: snow removal, rain streak and fog removal, and raindrop removal. Tab. 4 shows the comparison of our MIRAGE and other state-of-the-art methods. MIRAGE consistently outperforms existing methods across almost all datasets except the PSNR performance for RainDrop. The significant performance gains over multiple weather degradations demonstrate the effectiveness of MIRAGE in handling diverse weather-related degradations. Especially, 0.3 dB improvement on PSNR over Histoformer Histoformer_SunRGWC24 and 1.05 dB improvements over MPerceiver AiHZW024 .
Table 5: Zero-Shot Cross-Domain Underwater Image Enhancement Results.
Method PSNR (dB, $\uparrow$) SSIM ($\uparrow$)
SwinIR liang2021swinir 15.31 .740
NAFNet chu2022nafssr 15.42 .744
Restormer zamir2022restormer 15.46 .745
AirNet li2022all 15.46 .745
IDR zhang2023ingredient 15.58 .762
PromptIR potlapalli2023promptir 15.48 .748
MoCE-IR zamfir2025moce 15.91 .765
MIRAGE -S (Ours) 17.29 .773
Table 6: Complexity Analysis. FLOPs are computed on an image of size $224\times 224$ using a NVIDIA Tesla A100 (40G) GPU.
Method PSNR (dB, $\uparrow$) Memory ($\downarrow$) Params. ($\downarrow$) FLOPs ($\downarrow$)
AirNet li2022all 31.20 4829M 8.93M 238G
PromptIR potlapalli2023promptir 32.06 9830M 35.59M 132G
IDR zhang2023ingredient - 4905M 15.34M 98G
AdaIR cui2025adair - 9740M 28.79M 124G
MoCE-IR-S zamfir2025moce 32.51 4263M 11.48M 37G
MoCE-IR zamfir2025moce 32.73 6654M 25.35M 75G
MIRAGE -T (Ours) 32.77 3729M 6.21M 16G
MIRAGE -S (Ours) 32.91 4810M 9.68M 27G
Zero-Shot Cross-Domain (i.e., Underwater) Setting. We evaluate our method’s generalization under a challenging zero-shot setting using real-world underwater images. As shown in Tab. 6, MIRAGE -S achieves 17.29 dB and 0.773 SSIM, surpassing the best prior method MoCE-IR zamfir2025moce by a clear margin (+1.38 dB PSNR), while being more compact. Importantly, our model has never seen underwater data during training. This demonstrates that our adaptive modeling not only fits mixed degradations but also transfers robustly to unseen, conditions.
Efficiency Comparison. Tab. 6 presents a detailed comparison of PSNR, memory usage, parameter count, and FLOPs. Our Tiny model (MIRAGE -T) achieves the best efficiency-performance trade-off: with only 6.21M parameters and 16G FLOPs, it outperforms all prior methods, including larger models like PromptIR potlapalli2023promptir and MoCE-IR-S zamfir2025moce . Notably, MIRAGE -T surpasses MoCE-IR-S by +0.26 dB while requiring less than half the computational cost. Even our Small variant (MIRAGE -S) exceeds full MoCE-IR in both PSNR (+0.18 dB) and FLOPs (27G vs. 75G). These results validate that our design achieves strong restoration quality without sacrificing computational efficiency.
Visual Comparison. Across various restoration tasks, MIRAGE consistently produces sharper structures, richer textures, and fewer artifacts compared to existing methods. Fig. 1 demonstrates its strong generalization and ability to recover fine-grained details under diverse degradations, showing superior performance than others. More visual results are provided in our Supp. Mat.
5.2 Ablation Study.
We conduct ablation studies to assess the contribution of key components in MIRAGE , as summarized in Tab. 7. Starting from an attention-only baseline (32.23 dB, 19.89M), we progressively integrate each module while reducing overall complexity. Removing the dynamic convolution branch (w/o DynamicConv) causes a 0.56 dB drop, indicating its importance for local spatial modeling.
Table 7: Ablation Study of MIRAGE -T on 3 Degradation Setting.
Ablaton Parms. Results
PSNR (dB, $\uparrow$) SSIM($\downarrow$)
att-only (Ours) 19.89 M 32.23 (-0.54) .912
w/o DynamicConv 9.43 M 32.21 (-0.56) .911
w/o C-MLP 7.01 M 32.39 (-0.38) .913
w/o Fusion (i.e. Cat()-Only) 5.71 M 32.57 (-0.20) .914
w/o CL & SPD 5.80M 32.63 (-0.14) .916
w/o SPD 6.10M 32.53 (-0.24) .914
MIRAGE -T (Full) 6.21M 32.77 .919
The channel-wise MLP (w/o C-MLP) also plays a critical role, with a 0.38 dB performance loss. Replacing gated fusion with naive concatenation (w/o Fusion) leads to a further 0.20 dB drop, confirming that explicit feature integration is more effective. On the regularization side, removing contrastive learning (w/o CL & SPD) or only the SPD module degrades performance by 0.14 dB and 0.24 dB respectively, highlighting the benefit of manifold-aware cross-depth alignment. Overall, each component contributes to the final performance. Our full model achieves the best balance between accuracy and efficiency with only 6.21M parameters and 32.77 dB PSNR.
6 Conclusion
In this paper, we present MIRAGE , a unified and lightweight framework for all-in-one image restoration. By decomposing features into attention-, convolution-, and MLP-based branches, MIRAGE captures global context, local textures, and channel-wise statistics in a complementary and efficient manner. To bridge the semantic gap across depths, we introduce a shallow-latent contrastive learning scheme that aligns early and latent representations via second-order feature statistics on the SPD manifold. This improves cross-stage consistency and enhances representation discriminability without additional inference cost. Extensive experiments across diverse degradations and cross-domain settings show that MIRAGE achieves state-of-the-art performance with minimal parameters and FLOPs. We believe its design offers a scalable path for lightweight, adaptive restoration, with potential extensions to video and multi-modal scenarios. Please refer to our Supp. Mat. for additional discussions, implementation details, and visual results.
Appendix A Experimental Protocols
A.1 Datasets
3 Degradation Datasets. For both the All-in-One and single-task settings, we follow the evaluation protocols established in prior works li2022all ; potlapalli2023promptir ; zamfir2025moce , utilizing the following datasets: For image denoising in the single-task setting, we combine the BSD400 arbelaez2010contour and WED ma2016waterloo datasets, and corrupt the images with Gaussian noise at levels $\sigma\in\{15,25,50\}$. BSD400 contains 400 training images, while WED includes 4,744 images. We evaluate the denoising performance on BSD68 martin2001database and Urban100 huang2015single . For single-task deraining, we use Rain100L yang2020learning , which provides 200 clean/rainy image pairs for training and 100 pairs for testing. For single-task dehazing, we adopt the SOTS dataset li2018benchmarking , consisting of 72,135 training images and 500 testing images. Under the All-in-One setting, we train a unified model on the combined set of the aforementioned training datasets for 120 epochs and directly test it across all three restoration tasks.
5 Degradation Datasets. The 5-degradation setting is built upon the 3-degradation setting, with two additional tasks included: deblurring and low-light enhancement. For deblurring, we adopt the GoPro dataset nah2017deep , which contains 2,103 training images and 1,111 testing images. For low-light enhancement, we use the LOL-v1 dataset wei2018deep , consisting of 485 training images and 15 testing images. Note that for the denoising task under the 5-degradation setting, we report results using Gaussian noise with $\sigma=25$. The training takes 130 epochs.
Composited Degradation Datasets. Regarding the composite degradation setting, we use the CDD11 dataset guo2024onerestore . CDD11 consists of 1,183 training images for: (i) 4 kinds of single-degradation types: haze (H), low-light (L), rain (R), and snow (S); (ii) 5 kinds of double-degradation types: low-light + haze (l+h), low-light+rain (L+R), low-light + snow (L+S), haze + rain (H+R), and haze + snow (H+S). (iii) 2 kinds of Triple-degradation type: low-light + haze + rain (L+H+R), and low-light + haze + snow (L+H+S). We train our method for 150 epochs (fewer than 200 epochs than MoCE-IR zamfir2025moce ), and we keep all other settings unchanged.
Adverse Weather Removal Datasets. For the deweathering tasks, we follow the experimental setups used in TransWeather Transweather and WGWSNet ZhuWFYGDQH23 , evaluating the performance of our approach on multiple synthetic datasets. We assess the capability of MIRAGE across three challenging tasks: snow removal, rain streak and fog removal, and raindrop removal. The training set, referred to as “AllWeather”, is composed of images from the Snow100K liu2018desnownet , Raindrop qian2018attentive , and Outdoor-Rain li2019heavy datasets. For testing, we evaluate our model on the following subsets: Snow100K-S (16,611 images), Snow100K-L (16,801 images), Outdoor-Rain (750 images), and Raindrop (249 images). Same as Histoformer Histoformer_SunRGWC24 , we train MIRAGE on “AllWeather” with 300,000 iterations.
Zero-Shot Underwater Image Enhancement Dataset. For the zero-shot underwater image enhancement setting, we follow the evaluation protocol of DCPT hu2025universal by directly applying our model, trained under the 5-degradation setting, on the UIEB dataset li2019underwater without any finetuning. UIEB consists of two subsets: 890 raw underwater images with corresponding high-quality reference images, and 60 challenging underwater images. We evaluate our zero-shot performance on the 890-image subset with available reference images.
A.2 Implementation Details
[…]
Model Scaling. We propose two scaled variants of our MIRAGE , namely Tiny (MIRAGE -T) and Small (MIRAGE -S). As detailed in Tab. A, these variants differ in terms of the number of MDAB blocks across scales, the input embedding dimension, the FFN expansion factor, and the number of refinement blocks.
Table A: The details our the tiny and small version of our MIRAGE . FLOPs are computed on an image of size 224 × 224 using a NVIDIA Tesla A100 (40G) GPU.
     MIRAGE -T      MIRAGE -S
     The Number of the MDAB crosses 4 scales      [3, 5, 5, 7]      [3, 5, 5, 7]
     The Input Embedding Dimension      24      30
     The FFN Expansion Factor      2      2
     The Number of the Refinement Blocks      2      3
     Params. ($\downarrow$)      6.21M      9.68 M
     FLOPs ($\downarrow$)      16 G      27 G
A.3 Optimization Objectives
The overall optimization objective of our approach is defined as:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{1}}+\lambda_{fre}\times\mathcal{L}_{\text{Fourier}}+\lambda_{ctrs}\times\mathcal{L}_{\text{SPD}}.$ (6)
Here, $\mathcal{L}_{\text{Fourier}}$ denotes the real-valued Fourier loss computed between the restored image and the ground-truth image, and $\mathcal{L}_{\text{SPD}}$ represents our proposed contrastive learning objective in the SPD (Symmetric Positive Definite) space.
Specifically, we adopt an $\ell_{1}$ loss that adopted in IR tasks potlapalli2023promptir ; zamfir2025moce ; li2022all ; cui2025adair ; ren2024sharing , defined as $\mathcal{L}_{\text{1}}=|\hat{x}-x|_{1}$, to enforce pixel-wise similarity between the restored image $\hat{x}$ and the ground-truth image $x$. $\mathcal{L}_{\text{Fourier}}$, as utilized in MoCE-IR zamfir2025moce ; cui2025adair , to enhance frequency-domain consistency, the real-valued Fourier loss, is defined as:
$\mathcal{L}_{\text{Fourier}}=\left\|\mathcal{F}_{\text{real}}(\hat{x})-\mathcal{F}_{\text{real}}(x)\right\|_{1}+\left\|\mathcal{F}_{\text{imag}}(\hat{x})-\mathcal{F}_{\text{imag}}(x)\right\|_{1},$ (7)
[…]
Appendix B More Method Details & Supplementary Experiments
B.1 1 Deg. Comparison
Table B: Comparison to state-of-the-art for single degradations. PSNR (dB, $\uparrow$) and SSIM ($\uparrow$) metrics are reported on the full RGB images. Best performance is highlighted. Our method excels over prior works.
(a) Dehazing
Method Params. SOTS
DehazeNet cai2016dehazenet - 22.46 .851
MSCNN ren2016single - 22.06 .908
AODNet li2017aod - 20.29 .877
EPDN  qu2019enhanced - 22.57 .863
FDGAN dong2020fdgan - 23.15 .921
AirNet li2022all 9M 23.18 .900
PromptIR potlapalli2023promptir 36M 31.31 .973
MIRAGE (Ours) 6M 31.46 .977
MIRAGE (Ours) 10M 31.53 .980
(b) Deraining
Method Params. Rain100L
DIDMDN zhang2018density - 23.79 .773
UMR yasarla2019uncertainty - 32.39 .921
SIRR wei2019semi - 32.37 .926
MSPFN jiang2020multi - 33.50 .948
LPNet gao2019dynamic - 23.15 .921
AirNet li2022all 9M 34.90 .977
PromptIR potlapalli2023promptir 36M 37.04 .979
MIRAGE (ours) 6M 37.47 .980
MIRAGE (Ours) 10M 38.01 .982
(c) Denoising on BSD68
Method Params. $\sigma$=15 $\sigma$=25 $\sigma$=50
DnCNN zhang2017beyond - 33.89 .930 31.23 .883 27.92 .789
IRCNN  zhang2017learning - 33.87 .929 31.18 .882 27.88 .790
FFDNet  zhang2018ffdnet - 33.87 .929 31.21 .882 27.96 .789
BRDNet tian2000brdnet - 34.10 .929 31.43 .885 28.16 .794
AirNet li2022all 9M 34.14 .936 31.48 .893 28.23 .806
PromptIR potlapalli2023promptir 36M 34.34 .938 31.71 .897 28.49 .813
PromptIR potlapalli2023promptir (Repdoduce) 36M 34.15 .934 31.50 .894 28.33 .807
MIRAGE (ours) 6M 34.23 .936 31.60 .896 28.36 .808
MIRAGE (Ours) 10M 34.25 .937 31.65 .898 28.38 .810
Single-Degradation. Tab. B presents the results of MIRAGE trained individually for dehazing, deraining, and denoising. Across all tasks, our method consistently surpasses previous state-of-the-art approaches, including PromptIR potlapalli2023promptir and its reproduced variant. Particularly in the denoising task on BSD68, our model achieves the best performance across all noise levels. Interestingly, while the single-task models perform competitively, they are slightly outperformed by the all-in-one version in dehazing and deraining. This observation suggests that different degradation types may exhibit shared structures, and that learning them jointly can lead to more generalizable representations.
Algorithm A DynamicDepthwiseConv
1:$\alpha\in\mathbb{R}^{B\times C\times H\times W}$triangleright Input feature map
2:$\alpha^{\prime}\in\mathbb{R}^{B\times C\times H\times W}$triangleright Output after dynamic depthwise conv
3:[Step 1] Generate Dynamic Kernel
4:$K\leftarrow\texttt{AdaptiveAvgPool2D}(\alpha)$triangleright Global context pooling
5:$K\leftarrow\texttt{Conv2D}(K,\ 1\times 1,\ \text{out\_ch}=C)$triangleright Linear projection
6:$K\leftarrow\texttt{GELU}(K)$triangleright Non-linear activation
7:$K\leftarrow\texttt{Conv2D}(K,\ 1\times 1,\ \text{out\_ch}=C\cdot k^{2})$triangleright Generate kernel weights
8:$K\leftarrow\texttt{Reshape}(K,\ [B\cdot C,\ 1,\ k,\ k])$triangleright Form depthwise filters
9:[Step 2] Apply Depthwise Convolution
10:$\alpha_{\text{flat}}\leftarrow\texttt{Reshape}(\alpha,\ [1,\ B\cdot C,\ H,\ W])$triangleright Prepare for grouped conv
[…]
12:$\alpha^{\prime}\leftarrow\texttt{Reshape}(\alpha^{\prime}_{\text{flat}},\ [B,\ C,\ H,\ W])$triangleright Reshape back to original shape
13:return $\alpha^{\prime}$
Figure A: The illustration of different designs of the proposed MDAB.
B.2 Details of the Design for the proposed Mixed Backbone.
To investigate the effectiveness of combining MLP, convolution, and attention mechanisms, we conducted an extensive design-level ablation study. The quantitative results are presented in Tab. 7 of the main manuscript. Here, we provide detailed visual illustrations of each design in Fig. A.
C-MLP. To strengthen channel-wise representation, we introduce a Channel-wise MLP module, denoted as $\operatorname{C\text{-}MLP()}$. Given the input feature map $F_{\text{in}}^{\text{mlp}}\in\mathbb{R}^{B\times C\times H\times W}$, we first flatten the spatial dimensions to obtain a sequence $F_{\text{in}}^{\text{mlp}}\in\mathbb{R}^{B\times C\times L}$, where $L=H\times W$. The C-MLP is implemented using two 1D convolutional layers with a GELU activation in between. The GELU function introduces non-linearity, enabling the model to learn more complex and expressive channel-wise transformations. After processing, the output is reshaped back to the original spatial format, yielding $F_{\text{out}}^{\text{mlp}}\in\mathbb{R}^{B\times C\times H\times W}$.
Dynamic Depthwise Convolution. The $\operatorname{DynamicDepthwiseConv()}$ module is designed to capture content-adaptive local structures and is employed in Alg.1 of our main manuscript. As detailed in Alg. A, the input feature $\alpha\in\mathbb{R}^{B\times C\times H\times W}$ is first passed through a global average pooling and two $1\times 1$ convolutions to generate a dynamic depthwise kernel for each channel and sample. The input is reshaped and convolved with the generated kernels using grouped convolution, enabling sample-specific spatial filtering. The resulting output $\alpha^{\prime}$ maintains the original resolution while embedding adaptive local information.
Algorithm B SPD Contrastive Learning Optimization Pseudocode
# $f_{en}$: encoder
# $f_{de}$: decoder
# $patch\_embedding$: shallow convolutional patch embedding
# $refinement\_conv$: the refinement block and the final convolution
# $spd$: compute SPD feature
for ${x}$ in loader: # load a minibatch x with n samples
$F_{\text{shallow}}$ = $patch\_embedding$(${x}$) # Convolutional Patch Embedding
$F_{\text{latent}}$ = $f_{en}$($F_{\text{shallow}}$)
$\mathbf{C}_{s}$, $\mathbf{C}_{l}$= $spd$($F_{\text{shallow}}$), $spd$($F_{\text{latent}}$) # Compurte SPD (Symmetric Positive Definite) manifold features
$\mathbf{z}_{s}$, $\mathbf{z}_{l}$ = $proj\_norm$($\mathbf{C}_{s}$), $proj\_norm$($\mathbf{C}_{l}$) # Projection and normalize
$F_{\text{recon}}$ = $f_{de}$($F_{\text{latent}}$)
$\hat{x}$ = $refinement\_conv$($F_{\text{recon}}$)
L = $\mathcal{L}_{\text{1}}$(${x}$, $\hat{x}$) + $\lambda_{fre}\times$$\mathcal{L}_{\text{Fourier}}$ (${x}$, $\hat{x}$) + $\lambda_{ctrs}\times$$\mathcal{L}_{\text{SPD}}$($\mathbf{z}_{s}$, $\mathbf{z}_{l}$) # total loss
L.backward() # back-propagate
update($f_{en}$, $f_{de}$, $patch\_embedding$, $refinement\_conv$) # SGD update
def $\mathcal{L}_{\text{Fourier}}$(a, b): # Real-valued Fourier loss
␣␣␣␣RefertoEq.2ofourSupplementaryMaterails.
return loss
def $\mathcal{L}_{\text{SPD}}$(a, b): # Real-valued Fourier loss
␣␣␣␣RefertoEq.5ofourmainmanuscript.
return loss
B.3 Details of the Proposed SPD Contrastive Learning.
As shown in Alg. B, our SPD-based contrastive learning aims to align shallow and latent representations by operating in the space of symmetric positive definite (SPD) matrices. Specifically, given the shallow features extracted from the convolutional patch embedding and the latent features produced by the encoder, we compute their second-order channel-wise statistics to obtain SPD representations. These matrices are then vectorized and projected through learnable MLP layers, followed by $\ell_{2}$ normalization to form contrastive embeddings. An InfoNCE-style loss is applied between the shallow and latent embeddings to encourage structural alignment across depth. This contrastive term complements the pixel-level and frequency-based objectives, promoting more discriminative and consistent feature learning without introducing any additional cost during inference. Importantly, by leveraging the geometry of second-order feature statistics, our approach implicitly regularizes the representation space, encouraging intra-instance compactness and inter-degradation separability. This geometrically grounded formulation bridges low-level signal priors with high-level contrastive learning, offering a principled and scalable solution to all-in-one image restoration.
B.4 Ablation Regarding the Optimization Objectives
Table C: Ablation Study of MIRAGE -T on 3 Degradation Setting.
Ablaton Parms. Results
PSNR (dB, $\uparrow$) SSIM($\downarrow$)
w/o CL & SPD 5.80M 32.63 (-0.14) .916
w/o SPD 6.10M 32.53 (-0.24) .914
w/o Fourier Loss 5.80M 32.70 (-0.07) .917
MIRAGE -T (Full) 6.21M 32.77 .919
Tab. C shows that replacing SPD-based contrastive learning with a standard Euclidean-space contrastive loss (w/o SPD) results in a clear performance drop, demonstrating the advantage of modeling second-order channel correlations on the SPD manifold rather than relying solely on first-order vector similarities. When the entire contrastive module is removed (w/o CL & SPD), performance degrades even further, indicating that aligning shallow and deep features is essential for effective representation learning. Moreover, removing the Fourier loss (w/o Fourier Loss) slightly reduces performance, suggesting that frequency-domain supervision provides additional benefits. Overall, the full model achieves the best results, confirming the effectiveness of jointly optimizing spatial, frequency, and SPD-manifold-based structural consistency. Note that throughout all the experiments, we set $\lambda_{ctrs}$ = 0.05 and $\lambda_{ctrs}$=0.1.
B.5 Shallow-Latent Feature Similarity
Besides the channel-wise similarity comparison provided in our main manuscript for denoising. We also find consistent findings in other degradation, i.e., raining and hazing. The corresponding channel-wise similarity across scales is provided in Fig. B.
Figure B: The cross-sclae channel-wise similarity matrix visualization for Denoising, Deraining, and Dehazing.
These observations reveal several important trends: (i) Despite the diversity of degradation types, a consistent pattern emerges across scales. Specifically, from the first to the fourth scale, the overall channel-wise similarity indicates substantial redundancy among feature channels. After channel reduction, the latent features become more decorrelated, which validates the rationale for applying contrastive learning between the latent and shallow (i.e., first-scale) features. (ii) Different degradation types exhibit varying degrees of channel redundancy. As illustrated in Fig. B, hazy images tend to produce more inherently independent features, whereas rain-degraded inputs show strong channel-wise redundancy even in the latent space. This suggests that degradations like haze may benefit from larger embedding dimensions to capture more expressive representations, while simpler degradations (e.g., rain) can achieve effective restoration with smaller embedding sizes due to their inherently redundant structure.
These insights open up new directions for adaptive and degradation-aware model design in future research. Notably, this trend is not limited to the three representative samples shown; we observe similar patterns consistently across the dataset in a statistical sense. We plan to conduct a more comprehensive and quantitative investigation of this phenomenon in future work.
Appendix C Additional Visual Results.
C.1 3 Degradation
Figure C: Visual comparison of MIRAGE with state-of-the-art methods considering three degradations. Zoom in for a better view.
Fig. C presents qualitative comparisons on representative cases of denoising, deraining, and dehazing, benchmarked against recent state-of-the-art methods. The proposed MIRAGE consistently yields more visually faithful restorations, characterized by enhanced structural integrity, finer texture details, and reduced artifacts. These results underscore the effectiveness of our unified framework in handling diverse degradation types while preserving high-frequency information and geometric consistency.
C.2 5 Degradation
Figure D: Visual comparison of MIRAGE with state-of-the-art methods considering low-light degradation. Zoom in for a better view.
For the 5-degradation setting, we provide visual comparisons for the low-light enhancement task in Fig. D. As illustrated, the proposed MIRAGE produces noticeably cleaner outputs with improved luminance restoration and better color consistency compared to MoCE-IRzamfir2025moce , demonstrating its robustness under challenging illumination conditions.
C.3 Composited Degradation
Figure E: Visual comparison of MIRAGE with state-of-the-art methods considering composited degradation (Low-light + Haze + Snow). Zoom in for a better view.
Figure F: Visual comparison of MIRAGE with state-of-the-art methods considering composited degradation (Low-light + Haze + Rain). Zoom in for a better view.
Fig. E and Fig. F present visual comparisons under more challenging composite degradations, namely low-light + haze + snow and low-light + haze + rain, respectively. As observed, our method reconstructs significantly more scene details and preserves structural consistency, whereas MoCE-IR zamfir2025moce tends to produce noticeable artifacts and over-smoothed regions under these complex conditions.
C.4 Zero-Shot Underwater Image Enhancement
Figure G: Visual results of MIRAGE for Underwater Image Enhancement. Zoom in for a better view.
Fig. G demonstrates that even when directly applied to unseen underwater images, our method is able to effectively enhance visibility and contrast, producing results that are noticeably clearer than the raw input and visually closer to the reference images. This qualitative evidence further validates the strong generalization ability of the proposed framework to unseen domains.
Appendix D Limitations and Future Work
While the proposed MIRAGE achieves new state-of-the-art performance on most all-in-one image restoration benchmarks, we observe that its deblurring performance still lags slightly behind MoCE-IR zamfir2025moce . We attribute this to the relatively compact model size of our current design, which favors efficiency over aggressive capacity. To address this, future work will explore scaling up the model size to be on par with larger architectures such as PromptIR potlapalli2023promptir , MoCE-IR zamfir2025moce , and AdaIR cui2025adair , aiming to further boost performance while maintaining the architectural elegance and efficiency of our design. Moreover, our current SPD-based contrastive learning leverages a conventional InfoNCE loss in Euclidean space after projecting SPD features. While effective, it does not fully exploit the intrinsic geometry of the SPD manifold. As part of future efforts, we plan to investigate geodesic-based contrastive formulations and Riemannian-aware optimization strategies, which may offer a more principled and theoretically grounded way to align structured representations across semantic scales.
Appendix E Broader Impact
Image restoration (IR) is a foundational task with widespread applications across photography, remote sensing, surveillance, autonomous driving, medical imaging, and scientific visualization. By proposing a unified and efficient framework capable of handling diverse degradation types with minimal computational cost, our work has the potential to benefit a wide range of real-world scenarios where image quality is compromised due to environmental or hardware constraints. The lightweight design of MIRAGE enables deployment on resource-limited edge devices, such as mobile phones, drones, or embedded cameras in IoT systems. This democratizes access to high-quality image restoration, which may positively impact users in low-resource settings or in critical applications like emergency response and environmental monitoring. From a research perspective, our modular design and SPD-based contrastive formulation may inspire further work on principled representation learning for restoration and beyond, encouraging more geometrically-aware approaches in low-level vision. While our method is intended for general-purpose image enhancement, we acknowledge that improved image restoration techniques can also be misused for deceptive media editing or surveillance applications that raise privacy concerns. We encourage future practitioners to adopt this technology responsibly and in alignment with ethical standards. Our code and models will be released with appropriate licenses and documentation to promote transparency and responsible use.